OpenStudy (anonymous):

how do I find the x and the y? 6x+5y=39 3x+5y=27

OpenStudy (anonymous):

6x + 5y = 39 -3x - 5y = -27 3x = 12 x = 4 can you find the y?

OpenStudy (anonymous):

to be honest i don't remember the name of this method, but if you see two of the same variables having the same coefficient. you can switch the signs of one of the entire equations. then you can add the two equations together leaving you with one variable. lastly you can substitute the variable you solved for.

OpenStudy (anonymous):

thankz!!!

OpenStudy (anonymous):

No problem, do you have any questions or did you figure it all out?

OpenStudy (anonymous):

na I think I got it ha thankz though

OpenStudy (triciaal):

linear combination aka elimination method

OpenStudy (anonymous):

thnx triciaal

OpenStudy (anonymous):

elimination method!! thank you :D that was gonna kill me